OpenStudy (jkbo):

PLEASE HELP write the equation of each parabola with its vertex at the origin directrix at y= -8

OpenStudy (mertsj):

If the directrix is y = -8 and the vertex is (0,0) then p = 8 x^2=4py Use that.

OpenStudy (jkbo):

I still dont undersatand @Mertsj

OpenStudy (mertsj):

Ummmmmm how about you replace p with 8???? Is that too tough?

OpenStudy (jkbo):

so it would be x^2 = 4(8) right?

OpenStudy (mertsj):